OpenStudy (anonymous):

Suppose that the value of cos x is (1/3). Without your calculator determine the exact value of cos(2x).

OpenStudy (anonymous):

thats easy as well

OpenStudy (anonymous):

double angle expansion

OpenStudy (anonymous):

cos2x= cos^2 x - sin^2 x = cos^2 x - ( 1-cos^2 x ) = 2cos^2(x) -1

OpenStudy (anonymous):

\[\cos2x=2\cos ^{2}x-1\]